### v1

# DFabric: Scaling Out Data Parallel Applications with CXL-Ethernet Hybrid Interconnects

Abstract.
Emerging interconnects, such as CXL and NVLink, have been integrated into the intra-host topology to scale more accelerators and facilitate efficient communication between them, such as GPUs. To keep pace with the accelerator’s growing computing throughput, the interconnect has seen substantial enhancement in link bandwidth, e.g., 256GBps for CXL 3.0 links, which surpasses Ethernet and InfiniBand network links by an order of magnitude or more. Consequently, when data-intensive jobs, such as LLM training, scale across multiple hosts beyond the reach limit of the interconnect, the performance is significantly hindered by the limiting bandwidth of the network infrastructure. We address the problem by proposing DFabric, a two-tier interconnect architecture. First, DFabric disaggregates rack’s computing units with an interconnect fabric, i.e., CXL fabric, which scales at rack-level, so that they can enjoy intra-rack efficient interconnecting. Second, DFabric disaggregates NICs from hosts, and consolidates them to form a NIC pool with CXL fabric. By providing sufficient aggregated capacity comparable to interconnect bandwidth, the NIC pool bridges efficient communication across racks or beyond the reach limit of interconnect fabric. However, the local memory accessing becomes the bottleneck when enabling each host to efficiently utilize the NIC pool. To the end, DFabric builds a memory pool with sufficient bandwidth by disaggregating host local memory and adding more memory devices. We have implemented a prototype of DFabric that can run applications transparently. We validated its performance gain by running various microbenchmarks and compute-intensive applications such as DNN and graph.
1. Introduction
Data-intensive applications running in today’s production clouds, such as graph processing (SIGMOD10Pregel, ), data analytics (Commun08mapreduce, ), and deep neural network (DNN) training (Neurocomputing17DNN, ) often operate with Bulk Synchronous Parallel (BSP) (Commun90BSP, ) or MapReduce paradigms, which requires iterative execution of two key stages: data computation across multiple computing units (e.g., forward and backward propagation in DNN training), and communication between these units (e.g., data shuffling in MapReduce). Such applications are compute-intensive and would be very time-consuming. For example, training a BERT model on a single TPU takes over 1.5 months (arXiv18bert, ).
To accelerate the process, emerging interconnects, such as CXL (Compute Express Link) (CXLSpec, ) and NVLink (Micro17NVLink, ), have been integrated into the intra-host interconnect topology to scale a host with an increased number of computing nodes (e.g., accelerator, GPU, TPU (ISCA23tpu, )) and to facilitate efficient communication between them, for example, by supporting high bandwidth direct point-to-point communication. To keep pace with the evolving computing throughput, the interconnect has experienced a substantial improvement in link bandwidth as shown in Table 1. The link bandwidth of those new interconnects exceeds that of conventional network links by an order of magnitude or more, and this gap is expected to expand rapidly (SIGMOD20pump, ).
Table 1. The latency and bandwidth of mainstream devices.
Memory Interconnect Network
Device DDR5 GDDR6 CXL 3.0 NVLink 4.0 InfiniBand / Ethernet
Latency ¡80ns ¡220ns ¡400ns (MICRO23sun, ) ¡8us (TPDS20Li, ) ¿2us (EuroSys23DiLOS, )
B.W. 50 400 120 900 25/50
[…]
However, due to the distance limit of interconnect bus (existing CXL is up to 2-m maximum distance (SC22CXL3, )), they have limited scale, e.g., rack level (TACO24Rcmp, ). Thereby, the common practice in enterprise or public clouds is with emerging fast interconnects (e.g., NVLink, CXL) within hosts and network interconnects (e.g., Ethernet) between hosts (arXiv20scalable, ). Due to the huge gap between interconnect bandwidth and network bandwidth, as will show in §2, the data-parallel job, such as Allreduce, across multiple hosts or racks beyond the reach limit of the interconnect, its communication efficiency is inevitably hindered by the slow network link between hosts.
The intuitive approach to addressing the gap is to build a NIC pool with its aggregated capacity larger than that of the interconnect, intended for efficient communications beyond the interconnect’s reach limit. However, building a separate pool for each node’s local interconnect is cost-prohibitive due to the significant bandwidth disparity. For example, to match the bandwidth of a CXL 3.0 link, it would be necessary to install more than ten $200$ Gbps NICs per host. Instead, we propose connecting hosts and computing nodes using fast interconnects at a larger scale, i.e., rack level, thus there are sufficient number of existing NICs from hosts to be leveraged to form a NIC pool for communications across racks.
Specifically, we propose DFabric, a two-tier interconnect architecture. For intra-rack interconnect, DFabric utilizes CXL fabric, a recent interconnect fabric introduced in CXL 3.0, capable of scaling to thousands of different types of nodes like processors, memory devices, accelerators, and NICs. The fabric extends intra-host interconnects to the rack level. Communications across any pair of computing nodes (CN) across hosts is ensured by the fabric’s high bi-sectional bandwidth. For inter-rack level interconnect, DFabric also disaggregates existing NICs from individual hosts using the CXL fabric, and aggregates them into a NIC pool. The NIC pool is accessed by any node within the rack and tends to offer a sufficient aggregate bandwidth to transfer across-rack traffic load. Given the insight that NICs in conventional datacenter racks are often underutilized due to the on-off traffic pattern, as measured in  (NSDI15GRIN, ), and the contention at the NIC pool can be minimized by using some data parallelism approaches that exploit traffic locality, e.g., hierarchical Allreduce, or adopting contention-free communication patterns, e.g., ring Allreduce, by expanding the pool capacity with additional NICs, DFabric achieves the goal of providing a cost-effective, efficient, and flexible means to bridge the bandwidth gap in inter-rack communications. We will show in Figure 2 that DFabric achieves the optimal communication efficiency by simply forming a NIC pool with existing NICs, when running ring Allreduce, where only one node is communicating with the pool.
However, the memory bandwidth (The large NIC pool may push bottleneck to the Integrated IO controller (Anuj16ATC, ; PicNIC19SIGCOMM, ; Rolf18SIGCOMM, ; Midhul24SIGCOMM, ; Hostping23NSDI, ) and limited number of memory channels located in each host.) may pose a bottleneck that hinders the full utilization of the NIC pool’s capacity. This is because the bandwidth of local memory, i.e., system-integrated memory, or a specific memory device like GDDR, can be less than the capacity of the NIC pool. Consequently, when performing direct memory access (DMA) from all NICs, the memory bandwidth becomes the constraining factor, reducing the achievable throughput of the NIC pool. As CXL fabric in CXL 3.0 (CXLSpec, ) allows devices to participate in host processor’s coherence domain while accessing memory, we disaggregate the local memory and additional memory devices a rack using CXL fabric, and map them into a single memory address space to build a shared memory pool that can be accessed coherently by any device. Thus, the network traffic received from the NIC pool can be written into multiple memory devices with the aggregated memory bandwidth larger than the NIC pool’s capacity. Similarly, NIC pool reads the data using DMA from multiple memory devices when sending the traffic via the pool. By complementing the NIC pool with a shared memory pool, DFabric minimizes the across-rack communication period, and the computing nodes can directly access the memory pool with load/store instructions (CXL.mem) without moving the data to the local, which unleashes the hardware computing throughput.
Besides, there are admittedly several other critical challenges in realizing DFabric, for example, CXL.mem load/store is synchronous, and cacheline-based instructions, which is inefficient for accessing a large memory area compared to DMA, how to seamlessly use pass-by-reference semantic provided by interconnects for intra-rack communication, given the applications are developed using the socket programming model, etc.(more in §2). We address these challenges in design and implemented a full functional DFabric prototype with four customized MPSoC FPGAs and one X86 server connected by optical fibers, instead of using simulator and NUMA nodes in prior works (ASPLOS24CCNIC, ; arXiv23case, ; HPCA24Salus, ; ASPLOS23TPP, ), which enabling to run different applications transparently on DFabric. We implement CXL 3.0-like memory protocols, such as CXL.mem and CXL.io, atop an academia lightweight conceptual hardware protocol stack, which are needed by memory pool and NIC pool. To evaluate DFabric, we run both microbenchmarks and real data-intensive applications, such as DNN training, graph processing, and show that DFabric reduces a geometric mean of 30.6% communication time compared to running them with a conventional ToR-based rack. DFabric also achieves 40.5% lower p99 tail latency running Redis.
2. Background and Motivations
Figure 1. Two kinds of rack’s architecture proposed by previous work (a) and the DFabric (b). We abstract the inter-rack network as peer-to-peer connections. The bottlenecks are tagged with lightning.
Figure 2. The communication time under different bottlenecks running the ring allreduce.
In this section, we overview the emerging interconnects, i.e., CXL and NVLink, and validate the communication inefficiency due to the huge gap in bandwidth between the interconnect link and conventional network link.
2.1. Emerging Interconnects
A broad spectrum of applications ranging from traditional high-performance computing (HPC) to data-intensive applications running in clouds like machine learning, data analytics, and graph applications have been significantly accelerated by exploiting massive processors and accelerators (e.g., GPUs, TPUs) in parallel. To utilize more computing nodes at scale and facilitate efficient communication between processors and accelerators, and among accelerators, a new class of fast interconnects is emerging, such as Nvidia NVLink (Micro17NVLink, ), AMD Infinity Fabric (ISSCC18zeppelin, ), and Intel CXL (CXLSpec, ), and integrated into the intra-host interconnect topology to provide unprecedented bandwidth and low latency.
Bandwidth and latency.    To incorporate the evolving computing throughput of accelerators, the interconnect bandwidth has seen a substantial improvement, In Table 1, we show that both CXL and NVLink bandwidths exceed CPU memory bandwidth, enabling accelerators to access CPU memory at full memory bandwidth. Besides, the latency of CXL and NVLink exhibit one order of magnitude higher than the memory latency, and CXL incurs an estimated $70$ns for each hop over a switch when scaling CXL using multi-level switching (ASPLOS23pond, ). DFabric incorporates these properties of the bandwidth and latency into its design.
Resource pooling.    CXL fabric introduced in CXL 3.0 scales up to $4096$ nodes with CXL fabric (CXLSpec, ). such as remote memory devices, accelerators (e.g., GPUs, FPGA), and hosts, without performance degradation. CXL 3.0 supports memory sharing by mapping the memory of nodes into a single physical address space, which can be accessed concurrently by hosts in the same coherency domain. Compared to CXL 3.0, NVLink has the following limitations: 1) its cache coherent extension in fact supports GPUs as CXL type 2 devices (CXL.cache). However, it does not support CXL type 3 devices that support memory pool; 2) NVLink scales to 256 nodes though connects only GPUs as nodes (Micro17NVLink, ); 3) NVLink as a propriety interconnect limits its usage beyond Nvidia’s hardware. Thus, DFabric relies on CXL fabric as the underlying interconnect due to its generality, scalability and open standard.
Pass-by-reference semantics.    CNs can access any memory location with CXL.mem load and store instructions, which avoids expensive memory allocations and copying. NVLink enables load/store primitives between GPUs only, which is similar to CXL.cache instructions. DFabric leverages pass-by-reference semantics for intra-rack communications.
2.2. Communication Bottleneck and Strawman Approach
DFabric is motivated by the communication inefficiency when running a large-scale compute-intensive workload (e.g., graph, DNN training) in enterprise or public cloud clusters. These clusters are commonly built using ToR-based rack architectures, as shown in Figure 1, using fast interconnects (e.g., NVLink) within hosts and slow network interconnects (e.g., Ethernet) between hosts. For example, the highest configuration on Tencent Cloud with a 16-host cluster connected with 25 Gbps Ethernet, where each host has 8 Nvidia V100 GPUs connected with NVLink 2.0 that delivers 300 GBps aggregated bandwidth (arXiv20Tecent, ). Theoretically, when running the job across multiple hosts or racks in parallel, its communication efficiency is limited by the slow network link bandwidth.
To validate it, we assume the bandwidth ratio between interconnects and network links is $10$. We run a Gloo-based ring Allreduce (gloo, ) over a dual ToR-based racks architecture (Figure 1) emulated using FPGAs (see §5 and §6 for the experiment setup). Figure 2 shows the bandwidth gap is too large to bridge by merely adding one or two NIC, while adding too many NIC per host is infeasible due to “scale tax” such as power consumption, expense and operational costs (sirius-sigcomm20, ; rdc-nsdi22, ; shuffle-infocom24, ).
Strawman Approach    By replacing the ToR-based network interconnect with a fast interconnect, strawman DFabric interconnects hosts with CXL fabric at a large scale (at least 10 host rack) and reuse their existing NICs to form a NIC pool with sufficient large aggregated bandwidth (10-NIC pool) across racks. We built a dual-rack strawman DFabric by using CXL-DoCE (see §5), and Figure 2 validates that the resultant completion time of DFabric is approaching the optimal one.
The NIC pool’s capacity aggregated by existing NICs may not exceed the bisection bandwidth of the inter-rack switches, thus contention may happen at the NIC pool if its capacity is saturated by traffic from multiple links. However, communication efficiency is also guaranteed by the fact that ring Allreduce is crafted to minimize the contention at any node, which is commonly used in large scale compute-intensive applications (JPDC09allreduce, ), thus only one node at a time (host 1 or host 2) communicates with the NIC pool via a single CXL link. This also applies to other data parallelism approaches, such as hierarchical allreduce, which exploits traffic locality. Furthermore, prior research has reported that the utilization of NICs and bandwidth in ToR-based racks is typically low (SIGCOMM09VL2, ; IMC09Kandula, ; SIGCOMM10Benson, ), a finding that supports the efficiency of the NIC pool. As a final resort to address contention issues, the system allows for the flexible addition of additional NICs.
2.3. Challenges
Although the high level idea of DFabric is simple, there are admittedly several critical challenges to make DFabric work in various environments.
C1: Memory bandwidth can be a bottleneck.    The memory bandwidth may pose a bottleneck that hinders the full utilization of the NIC pool’s capacity. This is because that 1) the interconnect link bandwidth can be larger than the memory access bandwidth, and 2) the resultant NIC pool capacity can be larger than the memory bandwidth by incorporating concurrent transfers via multiple interconnect links. Consequently, when performing direct memory access (DMA) from all NICs, regardless sending or receiving, the memory bandwidth becomes the constraining factor, reducing the achievable throughput of the NIC pool. Figure 2 shows that DFabric’s performance is degraded when we reducing the memory access bandwidth intentionally.
To address C1, we leverage resource pooling in CXL 3.0 (§2.1) to disaggregate both local and remote memory devices using the CXL fabric. These devices are mapped into a unified memory address space, creating a shared memory pool accessible in a coherent manner by any CNs. Consequently, network traffic incoming from the NIC pool can be distributed across various memory devices, leveraging the aggregated memory bandwidth that surpasses the NIC pool’s capacity. Likewise, NIC pool performs direct memory access (DMA) reads from multiple memory devices to retrieve data for transmission through the pool.
C2: far memory access exhibits a longer latency.    Table 1 shows that the memory access with CXL memory protocols, i.e., CXL.mem, demonstrates latency approximately an order of magnitude higher than that of accessing local memory. The load and store of CXL.mem are synchronous, cacheline-based instructions with restricted concurrency, e.g., a maximum of 64 instructions (ISSCC22sapphire, ). This significantly impairs both intra- and inter-rack communication efficiencies, particularly when data is stored in remote memory (§4.6), as shown in Figure 2, DFabric’s performance is degraded to 2.1x. This is because hosts need to load packets from the memory pool. The inefficiency is pronounced for bulk transfer (footprint, ), compared to using DMA. To address C2, we introduce a DRAM cache in CXL ports, enabling caching in CXL.mem to effectively hide the latency (see §4.6).
C3: Compability.    The majority of DCN applications are developed using the socket programming model, which depends on the TCP/IP stack for intra-rack communication (§4.5). To address C3 and facilitate the use of pass-by-reference semantics, we introduce a kernel module that seamlessly translates socket system calls, such as SEND and RECV, into CXL.mem instructions such as load and store (see §4.5).
C4: NIC pool bandwidth sharing.    Cross-rack traffic traverses multiple paths by utilizing all available NICs. This raises the risk of out-of-order packet arrivals, potentially disrupting the in-order semantics of TCP. To address C4, DFabric relies on the memory pool to buffer out-of-order arrivals and let the host OS to sequence packets with Multipath TCP (MPTCP) (see §4.4).
Figure 3. An example architecture of the future datacenter. The resources within the DFabric are interconnected with CXL, while the racks are linked through Ethernet.
3. DFabric Overview
Architecture.    Figure 3 delineates the system architecture of DFabric. In DFabric’s architecture, traditional hosts are streamlined to CNs, each equipped with a processor or accelerator, such as a CPU or GPU, and integrated local memory. A segment of this local memory is disaggregated and, in conjunction with additional remote memory devices, constitutes a logical shared memory pool. The remaining local memory is designated as the private memory of the CN. This memory pool is integrated into a unified virtual address space, enabling consistent access through CXL.mem and CXL.io protocols. The NIC pool, which can incorporate a variety of NIC types, further enhances the system’s flexibility. A special CN, low power processor unit (LPPU) (SIGCOMM23Cowbird, ), is dedicated for some bookkeeping tasks such as enumerating, registering, and managing NICs. As detailed in §4.2, the LPPU virtualizes the NIC pool by integrating it into the address space as a singular, “big” logical NIC, which exposes the pool to all CNs and enables packet-based scheduling for cross-rack traffic, optimizing traffic distribution and bandwidth allocation. All resource nodes, including CNs, remote memory devices, the LPPU, and NICs, are interconnected by CXL fabric. CNs access remote resource devices via CXL switching in the fabric, forming a logical switch node (SN).
Workflow.    Although the underlying data structure (lock-free ring buffer) is not new and has been employed before (LITE17SOSP, ; FaRM14NSDI, ; Rambda23HPCA, ), our abstraction unifies inter- and intra-rack communication using the same Socket programming model. For intra-rack communication, CNs are communicated with pass-by-reference semantics, which uses CXL.mem load/store to pass references without moving the data (§4.3). For inter-rack communication, NIC pool DMA transaction is split into two parts (§4.4). Every CN allocates multiple virtual TX/RX queue pairs in its local memory. At the sending rack, a CN writes the packet descriptor to a virtual TX queue in its local memory. For each CN, there is a specific ASIC at the port to poll the descriptors from those queues. When reading a valid descriptor, the ASIC writes the descriptor to the working queue of a NIC. All above reads and writes use CXL.mem load and store instructions. The mapping between NICs and virtual TX/RX queues is determined by LPPU. LPPU is also responsible for NIC scheduling (§4.2) and memory pool allocation (§4.1). In §4.4, we design a NIC pool scheduling policy on a subflow basis.
At the receiving rack, receive buffers are allocated from the memory pool with the aggregate memory bandwidth larger than the NIC pool capacity. Thus, packets can be DMAed to the memory pool at the full speed of the NIC pool. There is a dedicated ASIC to poll all the completion queues of the NIC pool. For every ready descriptor, the ASIC writes the descriptor to the corresponding virtual RX queue of the CN based on the destination IP of the descriptor, and then interrupts the CN.
4. Design Details
In this section, we present a detailed DFabric designs, especially for those addressing the above challenges.
Figure 4. The management machanisms on memory pool
4.1. Memory Pool
A unified addressed memory pool is the key to realize the pass-by-reference intra-rack communication and two-stage inter-rack communication. The memory pool is mapped to a single fabric address space (FAS) and accessed with CXL.mem load/store. We complement CXL fabric with management mechanisms on the memory pool such as organization, bootstrap, and allocation, which pave the foundation for intra- and inter-rack communications.
Organization.    LPPU organizes the shared memory pool as a series of coarse-grained Sections, each with a size equal to that of a huge page ($2$ MB). LPPU also groups $N$ consecutive Sections into a region, where $N$ is configurable, and N=512 in our case. Upon a Section is allocated to a CN, the Section can be further divided into Buffers by CN with the buffer size set based on the running application’s demand (with APIs in Table 2), e.g., the Section with 1 KB Buffer.
Bootstrap.    LPPU enumerates remote memory devices attached by CXL fabric and CNs’ local memory in the memory pool, and builds a mapping table that maps FAS to memory physical addresses. Then, every CN enumerates the FAS and private memory, so that it can access the pool with CXL.mem. Note that the CN’s private memory has a different address space, which is used for kernel, and caches program codes and hot data. We further use the private memory to implement DRAM cache, which is used to hide the non-uniformed latency of the memory pool from CNs (§4.6).
Memory allocation.    A daemon running on every CN is responsible for allocating Sections in advance and managing them. LPPU updates the mapping from the corresponding FAS to the Section’s physical address in the CN’s mapping table. Similar to the Linux buddy subsystem, the daemon applies multiple Sections, each having a different Buffer size, the minimum memory management unit, e.g., a Section with 2 KB Buffers. An application running in a CN will consult a daemon for user-level memory allocations when it calls alloc_shared_buffer. Specifically, the daemon will go through the following steps: 1) it chooses Sections with a desirable Buffer size based on the application semantics, 2) it chooses a Section based on the physical location of the Sections. This is because, from the CN’s perspective, the accessing latency profile of the memory pool is different based on the memory’s physical locations. e.g., we prefer local Sections to store virtual queues to minimize accessing latency.
Discussion.    DFabric relies on CXL port to translate the memory request FAS to the physical addresses and verify the permission with the CN’s mapping table, DFabric routes requests with the routing algorithm in CXL fabric (CXLSpec, ), e.g., Port-based.
Figure 5. Data structure and metadata used in the interactions with NIC pool and memory pool.
4.2. NIC Pool
DFabric decouples NICs from hosts to form a NIC pool and interconnects them via CXL fabric. CXL fabric provides CNs with a single NIC abstraction. LPPU is responsible for the control plane of the NIC pool including configuring NICs and NICs scheduling. To schedule the NIC pool capacity efficiently and flexibly, LPPU schedules NICs on a packet-basis based on their working status, e.g., NIC’s working queue depth.
The NIC pool and its metadata are shown in Figure 5. The metadata is used for both intra- and inter-rack communications. LPPU abstracts NIC pool as a single NIC to each CN via a set of virtual queues (virt_queue) that are stored in the memory pool. virt_queue includes RX queues (RxQ) and TX queues (TxQ). LPPU maintains every NIC’s working queue (physical_queue) in its private memory, and uses them for NICs scheduling (see §4.3). CNs and LPPU are responsible for initiating virt_queues and phy_queues respectively at bootstrap.
ASIC is responsible for the data plane between CNs in intra-rack communication (§4.3), and CNs and NIC pool in inter-rack communications (§4.4). For example, ASIC interacts with the NICs by storing and loading descriptors to and from physical_queues.
4.3. Intra-rack Communication
DFabric’s intra-rack communication is based on pass-by-reference (pointer) semantics without moving data. To identify the destination of the references, every CN has a unique ID (e.g., IP). The communication semantic is to transfer the reference addressing the data in the shared memory from CN $ID_{src}$ to CN $ID_{dst}$, so that CN $ID_{dst}$ can load/store it. To transfer a reference, DFabric stores a transaction descriptor into a virt_queue, where every entry contains a transaction’s $ID_{dst}$ and $ID_{src}$, a reference, and the length of the data it references. As shown in Figure 5 the intra-rack communication can be summarized into the following steps:
Data preparation.    $ID_{src}$’s application or runtime directly updates the Buffer content via load/store to the address ➀. CXL port translates the instruction’s FA to physical address, and encapsulates it in the format of CXL.mem transaction layer packet (TLP).
TxQ operations.    $ID_{src}$ stores the Buffer’s descriptor to the head entry of its TxQ, where TxQ is implemented as a circular buffer. The ASIC of $ID_{src}$ polls all TxQs and obtains a new descriptor by checking the tail of a TxQ ➁.
Split-transaction in ASIC.    The ASIC checks whether the descriptor’s destination is the NIC pool or another CN. For the former one, inter-rack communication is instantiated (see §4.4). Otherwise, ASIC stores the Buffer’s descriptor to the head entry of RxQ of $ID_{dst}$ and writes to its interrupt register.
RxQ operations.    $ID_{dst}$’s runtime loads the tail entry of RxQ, and informs the application with the data reference ➃.
Data buffer deallocation.    Once the application or runtime no longer uses Buffer, it will free the Buffer’s address to the daemon. If there are enough Buffers, the daemon will further free the Buffer to LPPU.
The proposed pass-by-reference mechanism leads to zero data copy, and works transparently with the application using the socket programming model (see §4.5). We can also apply the pass-by-reference mechanism to DPDK (dpdk, ) and RDMA verb similarly.
4.4. Inter-rack communication
Despite NIC pool allows a CN to leverage multiple NICs to achieve a higher throughput. we face two challenges. 1) As the working status of every NIC’s physical_queue and the path condition behind it vary over time, it is important to design a flexible and efficient NICs scheduling policy that can operate on a finer granularity, e.g., packet basis, while fully utilizing the NIC pool’s capacity. 2) When a flow is scheduled with multiple NICs, it is hard to ensure packets arrive in-order as the path delay can differ. If the underlying transport enforces in-order arrivals, e.g., TCP, out-of-order arrivals will trigger the congestion control to reduce the sending throughput unnecessarily. Thus, it is also important for the NICs scheduling policy to minimize the out-of-order arrivals.
DFabric’s NICs scheduling.    CNs equally divide every flow into subflows and map each to a TxQ/RxQ pair. LPPU many-to-one maps TxQs to NICs based on the utilization of NICs, e.g., the NIC’s physical_queue depth (such as using hardware counters of NVIDIA RNICs through network adapter management tool NEO-Host (NVIDIANEO, )). Note that we assume that there exists no link down in the core network, and every sub-flow is mapped to a fixed path via ECMP when it traverses the network core. Thus, DFabric fully exploits the NIC pool only if CNs can generate a sufficient number of subflows, while ensuring in-order arrivals at the destination within a subflow. To this end, we exploit MPTCP-like indirection in the sending CN’s OS, which can open a sufficient number of TCP subflows, while resequencing TCP subflows into an in-order data stream at the receiving CN (mptcp_rfc, ).
For inter-rack communication, we assume every NIC has a three physical_queues; TX_phq, a RX_phq, and a completion_phq, as shown in Figure 5.
Packet transmission.    The ASIC of $ID_{src}$ polls all its TxQs in the memory pool. For every TxQ, LPPU has scheduled a NIC to send its Buffers referenced by the entries in the TxQ. The ASIC stores DMA descriptors in that NIC’s TX_phq ➄. The NIC’s DMA engine is then notified to collect data from the Buffers referenecd and perform network packetizing.
Packet receiving.    LPPU allocates multiple Buffers in the memory pool as pool’s receiving buffers in advance. The receiving buffers for NICs should have sufficient memory bandwidth than the throughput of the NIC pool. LPPU also fills the RX_phq of every NIC with DMA descriptors referencing receiving Buffers in advance. For each packet’s arrival, the NIC’s DMA engine directly writes the payload to the Buffer specified by the valid in-queue DMA descriptor. A completion notification is stored in the NIC’s completion_phq as soon as the DMA engine finishes a DMA operation ➅. A dedicated ASIC is responsible for polling these completion_phqs and informing the CNs of data arrival, by storing the CN’s interrupt register based on the packet’s IP and storing the references of the Buffers in the RxQ with CXL stores ➂. Eventually, the CN can load/store Buffers during its computation.
Table 2. memory pool related APIs
Name Input Output
build_shared_skb buffer reference, length sk_buf reference
kfree_shared_skb sk_buf reference Void
alloc_shared_buffer size, location buffer reference
4.5. Software Runtime
To enable applications to enjoy pass-by-reference semantics transparently, DFabric provides a set of APIs for TCP/IP stack, and a driver below the stack to orchestrate intra-rack communication.
The TCP/IP stack operates on the socket buffer object (sk_buf), which includes the data field that stores the reference to the actual data, and the data can be located at any memory device within the memory pool. For performance consideration, the TCP/IP stack uses the write-around cache policy for these buffers so that the data is directly written/updated in memory without bringing them to the cache first. Recall, the daemon in every CN is responsible for applying Sections from LPPU. Applications and runtimes can apply Buffers from the daemon, which can be used for user memory and kernel memory, respectively.
Socket buffer APIs.    To enable TCP/IP stack to read these Buffers in the shared memory pool, we define three APIs in Table 2. Recall, CN will be interrupted to read RxQs whenever new references are available. CN $ID_{dst}$ calls build_shared_skb to wrap the received references with sk_buf before passing it up to TCP/IP stack. Once TCP/IP no longer uses the buffer, it calls kfree_shared_skb to hand over the buffer to the daemon. CN $ID_{src}$ should call alloc_shared_buffer to apply for a buffer from the daemon before populating the buffer with the application’s data and TCP/IP headers. As the memory pool presents a non-uniform accessing latency, $ID_{src}$ is encouraged to pass the preference on Buffer locations based on the application semantics. For example, a Buffer located at the SN is preferred if several CNs will own the buffer alternately.
DFabric driver.    For configuring DFabric to orchestrate intra-DFabric communications, the driver exposes network device operators (ndo) to TCP/IP and spawns the daemon. The operators include 1) setting up unified memory address space at the bootup stage by registering contiguous FAS to CNs and local shared memory to the memory pool; 2) initiating intra-rack communication transactions once TCP/IP has prepared sk_buf; 3) handling interrupts when receiving communication transactions.
Figure 6. The architecture and working flow of the DRAM cache.
4.6. CXL-attached DRAM Cache.
CXL load/store is synchronous and fine granularity memory accessing, which limits the throughput of accessing Buffers allocated in remote memory devices, which exhibits locality on memory accessing pattern. For example, copying consecutive data from the kernel to the user mode may exhibit a high spatial-locality. As CXL and the CPU structure (MSHR) limit the maximum outstanding number of load/store instructions, reducing memory latency via caching or migration is the only way to improve the bandwidth.
Recent works (MACS23DaeMon, ; ASPLOS21Kona, ; ICCD23Morpheus, ) validated that CXL-attached accelerators can track memory accesses at the cache-line granularity, and we can upgrade or replace them without modifying the entire chassis. Compared to caching, page migration (ASPLOS23TPP, ; ASPLOS23pond, ; TACO24Rcmp, ) requires the runtime to identify hot pages, migrate them to local memory, and update the page table. we design to install a DRAM cache card in one of the CXL ports at all CNs.
By analyzing the TCP/IP stack, we summarize the three key points to guide the DRAM cache design. 1) Variable time interval between accessing headers and payloads. The runtime processes packet headers and data during the bottom half of the interrupt, and the recv syscall, respectively. Therefore, there will be a time interval between the two, in which the data can be evicted out of the cache. 2) Buffer release function is explicitly called. As runtime calls the buffer management APIs explicitly, it will flush the Buffer out of the cache once it calls the buffer release function, i.e., kfree_shared_skb. 3) Variable packet granularity. The packet size may be a few cache lines, such as the TCP SYN and ACK. Caching data with coarse granularity leads to a low cache utilization rate and prolongs the latency, thus the packets should be filtered before caching.
As shown in Figure 6, we use a multi-way DRAM cache to tackle the potential eviction caused by conflict during variable intervals. The metadata (Tags) is stored separately in the on-chip memory, so that one read can get the whole set’s metadata. The usage of DRAM cache can be summarized as follows. ➀ The driver configures the FAS of a region – a consecutive number of Sections, whose Buffers will be cached in DRAM cache, and the caching granularity, i.e., Buffer size. For the first access to a Buffer in the Region ➁, a miss is triggered and the Buffer is fetched from the memory pool via CXL.io ➂. Then, the DRAM cache fills valid metadata and evicts the victim Buffer. The following accesses to the Buffer ➁ will hit the local DRAM cache. When the runtime decides to free Buffers, kfree_shared_skb will be explicitly called and flush the DRAM cache.
Currently, DFabric uses the DRAM cache to improve the throughput of loading/storing a large memory segment. However, the cache design is general and independent from the runtime, thus applications can transparently use the DRAM cache by configuring their FAS of Regions.
Figure 7. The architecture of each port’s data plane accelerator.
4.7. ASIC Architecture
Each CXL port of SN has a data plane accelerator implemented using ASIC or FPGA (Figure 7), which is responsible for the intra- and inter-rack communication defined in Section 4.4 and 4.3. Despite functions defined by CXL Switch, such as forwarding requests among ports, address decoding, etc, the Coherence Controller handles the memory requests from other components and forwards the polled transaction descriptors to the Dispatcher. the Coherence Controller may instantiate an address translation cache to support applications’ virtual address (ATS, ; Kong23NSDI, ). The Dispatcher forwards descriptors to other ASICs or the Processing Unit based on which type of communication the descriptors define. The Processing Unit should be general-purpose so that it can translate transaction descriptors to specific descriptors of various NICs. Besides, the Processing Unit polls or read/write NICs’ physical_queue via the Coherence Controller, and translates completion notifications to transaction descriptors. SN may implement multiple Processing Units distributed within each ASIC, or a giant Processing Unit shared by all ASICs. SN’s private memory can be a separate CXL memory device or RAM integrated with the Processing Unit. The Virt-Queue Manager contains the context of each virtual queue, including the head, tail pointers, and queue depth. It needs to poll the TxQs and store received transaction descriptors from other ASICs and the Processing Unit in the RxQs via the Coherence Controller.
5. Implementation
Considering no commercial products supporting CXL 3.0, we implement a proof-of-concept dual-rack system prototype as shown in Figure 8. We use this prototype to emulate a dual-DFabric architecture as well as dual-ToR-based racks architecture as the baseline (§6.1). In a rack, there are two customized MPSoC FPGA (mpsoc, ) as CNs. Each FPGA has a quad-core ARM CPU, two memory channels, and four optical fiber ports. The MPSoC exports the CPU memory bus to FPGA logic via HP/HPC ports, which can be used to implement a CXL-like load/store. We connect two FPGAs to a dual-port Intel 82599 NIC (82599, ) installed on an x86 server. We use that x86 server to emulate two SNs for DFabric, and two ToR switches for the baseline systems. We emulate NIC pool and inter-rack communication with a DPDK-based network emulator  (dpdk, )), running in the x86 server.
CXL-like protocol layer.    The key to emulating CXL fabric in CXL 3.0 is to externalize loads and stores. This allows applications to access remote memory devices within the memory pool transparently, enabling the seamless execution of legacy applications. To this end, we leveraged DoCE (ETCD17doce, ), a hardware protocol stack as the basis of our CXL-like protocol layer. DoCE directly encapsulates all ARM AMBA AXI on-chip interconnect signals within an Ethernet frame which can be delivered via standard Ethernet infrastructure. Instead of packing AXI signals into Ethernet packets with DoCE, we augmented DoCE with a CXL-like protocol layer and implemented them as a hardware module in the MPSoC FPGA (i.e., CXL-DoCE), which transforms AXI signals into relevant CXL transactions, such as MemRd, MemWr, Cmp, and MemData defined in CXL 3.0 specification (CXLSpec, ), before encapsulating CXL transactions into Ethernet packets. As shown in Figure 8, any memory transaction accessing remote memory from the CPU would go through the hardware, which transforms it into an Ethernet packet with a CXL-like transaction and sends it via an optical fiber port. Similarly, any received Ethernet packet from the port would go through the module, which transforms it into AXI signals to access the local memory if that packet contains CXL-like transactions. Note that an Ethernet packet can encapsulate transactions not limited to CXL-like as long as the module can recognize them, e.g., the module can interrupt the ARM CPU if the Ethernet packet carries an interrupt transaction.
Advantages.    Compared to previous works that leverage NUMA nodes to emulate CXL-like load/store, i.e., the load/store to the memory of the remote socket is emulated as CXL ones, our platform can adjust the latency of load and store and forward CXL-like transactions across multiple devices via Ethernet to emulate a CXL fabric, rather than being constrained by the two-node (socket) scenario (ASPLOS23pond, ; ASPLOS23TPP, ; ASPLOS24CCNIC, ). Compared to implementing DFabric in a simulator, our prototype can run real applications within a reasonable time. For example, full system mode Gem5 is thousands of times slower than the real-system  (arXiv20gem5, ), which makes it unable to simulate full application execution.
6. Evaluations
Figure 8. The dual-rack simulating platform with 4-FPGAs array (a) is the customized MPSoC FPGA marked with key hardware components and the software emulator sunning on the X86 server. (b) is the RTT latency of accessing memory with/without added cycles, measured by pointer chasing from Lmbench (ATC96lmbench, ).
In this section, we evaluate DFabric’s performance gains compared to the baseline ToR-based rack architecture. Specifically, we address the following questions by running various experiments on our proof-of-concept prototypes. 1) How does DFabric perform when running legacy data-intensive applications? (§6.2) 2) How does DFabric perform regarding intra-rack latency and inter-rack bandwidth with varied parameters? (§6.3) 3) The breakdown of the contribution of the key designs to the DFabric’s performance gain. (§6.4)
Table 3. Key parameters of the two prototypes.
Parameter Value
DFabric
Memory Pool 3x16 GB; remote 6.5 us; local 650 ns;
SN 1GB private memory; 2+M NICs (Uplink)
CN 2GB DRAM Cache; 1 CXL port
Baseline
ToR 1 Uplink with B Mbps; 256 KB per port (Wei-nsdi16, ; Bai-CoNext16, ; buffer_sz, )
CN 1 NIC with B Mbps
Intra-rack Network 32.5 us; B Mbps;
Common
Inter-rack Network 32.5 us; $B\times\#\text{Uplink}$ Mbps;
ARM CPU 4-core A53; 1.2 GHz
x86 Server 64-core Xeon Gold 6130; 3.7 GHz
6.1. Experiments Setup
We set MTU to be 4 KB for both prototypes, and the 4KB Buffers belong to the Regions which are cacheable by the DRAM Cache. The configuration parameter values used in the experiments are summarized in Table 3.
Baseline rack architecture for comparison.    As shown in Figure 8, we have ported an example project (baseline, ) to our MPSoC FPGA, in which the ARM CPU uses one optical fiber port as a NIC. We also use a DPDK-based network emulator running on the x86 server to emulate ToR switches and the network links. We configure each network hop’s latency as the same ratio as RDMA compared to CXL memory (5:1) (EuroSys23DiLOS, ; MICRO23sun, ). We use DFabric and Baseline to denote the dual-DFabric prototype and the dual ToR-based racks prototype, respectively, for the later discussion.
Latency alignments.    To align the access latency ratio between local memory and CXL-attached remote memory, as shown in Figure 8, we add a hardware module on MPSoC FPGAs to add reconfigurable cycles to all memory accesses from ARM. According to the real measurements of those two latencies (MICRO23sun, ), we configure the latency ratio between the two to be 1:10 by adding $\lambda$ cycles to both memory accesses, such that $\frac{D_{local}+\lambda}{D_{cxl}+\lambda}=\frac{1}{10}$, where $D_{local}$ and $D_{cxl}$ denote the access latency of local memory and CXL-attached remote memory respectively. As shown in Figure 8, we increase the data array size until accessing it triggers a cache miss, e.g., the data array size from $256$ MB to $1024$ MB, and use the resulted $D_{local}$ and $D_{cxl}$ to compute $\lambda$, and find when $\lambda=120$, both access latencies meet the 1:10 ratio.
Bandwidth alignments.    To reflect the bandwidth gap between CXL and Ethernet links, we limit the NIC bandwidth with a configurable reducing factor $\theta$. Thus, the Ethernet link bandwidth in Baseline is set to $B=\frac{C}{\theta}$, where $C$ denotes the maximum achievable throughput of the CXL. Note that DFabric forms a NIC pool with both the NICs of the CNs and newly-added NICs, thus the capacity of the NIC pool is $\frac{C}{\theta}\times(N+M)$, where $N$ and $M$ denotes the number of CNs within a rack and the number of newly-added NICs, respectively.
Figure 9. Comparison between Baseline and DFabric on real-world applications. The X-coordinate is the configured bandwidth of each NIC (B) if not specified. Figure f, g, h is sampled with $B=\frac{C}{8}$.
6.2. Real Application Workloads
We ported the following real-world application frameworks to DFabric transparently: MapReduce (mrlite, ), GeminiGraph (OSDI16Gemini, ), PyTorch Distributed Data Parallel (DDP) (arXiv20pytorch, ), and Redis (carlson2013redis, ). To measure the metrics of synchronization stages shown in Figure 9, we use the pdump (dpdk, ) running on the X86 server to capture the transferred packets, which will then be analyzed by the WireShark (lamping2004wireshark, ). We increase the severity of network bottlenecks by decreasing the parameter B in Figure 9.
Graph Processing.    We run the PageRank and BFS using the LiveJournal online social network (snapnets, ). As shown in Figure 9 9, DFabric can reduce the communication time by an average of 59.5% in the worst case ($B=\frac{C}{8}$), and 32.1% in all cases. Each CN will finish iterating its vertices asynchronously, so that CNs can use the NIC pool exclusively during the synchronization stages. As saving packets to the memory pool alleviates issues like incast, the performance of DFabric at point M (NIC pool with $\frac{C}{4}$) is better than the Baseline at point N (2 NICs with $\frac{C}{4}$). We present the communication throughput sampled during one run of PageRank in Figure 9. There exist 12 bandwidth peaks as the same number with supersteps, and DFabric greatly reduces the communication time compared to Baseline.
Figure 10. The communication time measured running LLM model.
Neural Networks.    We train the residual neural network ResNet18 (arXiv15resnet, ) using the CIFAR-10 dataset (krizhevsky2009CIFAR, ), and the communication time is shown in Figure 9. ResNet18 has around 11M parameters synchronized among all CNs after training each batch. We use Gloo as the communication backend for PyTorch, which coordinates CNs as a ring. In this case, only one CN will use the NIC pool to transmit or receive packets among racks, so that the DFabric possesses higher inter-rack throughput than Baseline. DFabric reduces the communication time by 54.1% and an average of 27.1% in the worst and all cases respectively. The throughput sampled during one of the synchronization stages is shown in Figure 9, and the coexistence of the NIC pool and CXL leads to a larger throughput than a single NIC.
We train a large language model TinyStories (arXiv23TinyStories, ) with 1M parameters on the same software stack with ResNet18 as shown in Figure 10 10. CNs conduct ALLtoALL communication schemes for gradient synchronization. The high throughput of CXL and the memory pool alleviating the incast issue lead to an average of 34.7% reduction in communication time.
WordCount.    We run the MapReduce WordCount by configuring one CN to perform reducing tasks and the other three CNs to run mapping tasks. Three flows to the reducer during the reducing phase may cause a severe incast issue at the reducer in the Baseline. In contrast, DFabric absorbs the burst inter-rack burst traffic with a rack’s memory pool with sufficient capacity and bandwidth, thus achieving a higher inter-rack throughput without packet loss, while offloads intra-rack communication to CXL fabric which does not move data but references. As a result, DFabric reduces communication time by an average of 31.1% shown in Figure 9. The throughput sampled during one of the reducing phases is shown in Figure 9. After intra-rack traffic is finished, the throughput is stable at nearly 250Mbps for the last minutes which is limited by the NIC pool.
Redis Cluster.    We run Memtier_benchmark (memtier, ) on one CN as a client and organize the other three CNs to form a Redis Cluster. The client sends Get/Set requests to different servers depending on the distribution of the key-value pairs. We monitor the average and p99 latency of requests because the incast issue will lead to packet loss which prolongs the tail latency. In DFabric, there is zero packet loss within the rack as references are passed instead of values, and the memory pool will store the inter-rack packets in time. As shown in Figure 9, compared to Baseline, DFabric reduces p99 and average latency by an average of 40.5% and 22.8% respectively. However, DFabric presents higher p99 latency when no network bottleneck exists ($B=C$) caused by the memory pool’s long accessing latency.
Figure 11. The intra-rack transmission latency comparison between pass-by-reference (DFabric) and pass-by-value (Baseline) based TCP/IP. There is no constraint on the NIC’s bandwidth for Baseline ($B=C$).
6.3. Communication Efficiency
Intra-rack Latency.    As DFabric eliminates one memory copy by using pass-by-reference within the rack, in this section, we evaluate the latency reduction from DFabric’s pass-by-reference TCP/IP protocol stack. We run the simple TCP transaction latency test from lmbench (ATC96lmbench, ), which repeats transmitting one packet filled with dummy data in a ping-pong manner. The legacy recv system call will copy data into the user buffer no matter whether it will be used. Several techniques have been developed that try to alleviate the latency introduced by this copying, such as re-mapping, copy-on-write techniques (Solaris96ATC, ), and offloading TCP/IP processing onto the NIC (DOJO22HotChips, ; Kim23NSDI, ). Therefore, We remove the copying action in the recv system call for latency testing only. As shown in Figure 11, DFabric presents an average of 15.9% lower latency than Baseline. The DFabric’s latency slowly increases along with the message size as the send system call will copy data from the user buffer to the kernel sk_buf.
Figure 12. The average bandwidth comparison between different numbers of newly-added NICs (M) in the NIC pool. We run 4 kinds of communication patterns from Gloo.
Inter-rack Bandwidth.    One of the advantages of the NIC pool is freely scaling up more NICs, and we evaluate the number of NICs’ impact on the inter-rack bandwidth by increasing M. We build 4 communication scenarios using Gloo (gloo, ), which is a collective communications library atop the TCP/IP stack, on DFabric and monitor the average bandwidth of CN0. For example, Gather and Broadcast scenarios require CN0 to receive/send data from/to others, while the All-to-All scenario simultaneously conducts the former two scenarios on each CN. The Ring-Reduce scenario coordinates CNs as a ring and manipulates each CN’s data to send to and receive from the left and right neighbors respectively. As shown in Figure 12, the average bandwidth increases with the number of NICs, gradually reaching the limit as the bottleneck is shifted to the CN’s processing rate. Because the CN sends and receives data simultaneously in All-to-All and Ring-Reduce scenarios, DFabric achieves lower bandwidth than other scenarios.
6.4. Deep Dive
Performance breakdown.    Although the high-level idea of DFabric is simple, we introduce several key designs to address the challenges in building an efficient DFabric, such as tcp small queue related to pass-by-reference TCP/IP stack, concurrent TxQs loads on SN’s ASIC design, DRAM cache. To figure out the performance contribution of the above designs to the DFabric efficiency, we compare the throughput reported by iPerf of the DFabric with a specific design disabled to the full-functional DFabric.
Table 4. The throughput breakdown of the DFabric.
Disable opt. w/o TCP small queue SN concurrently LD TxQ DRAM Cache
ratio 0.50 0.75 0.17
As shown in Table 4, the resulting throughput is normalized to full-functional DFabric’s throughput. Disabling the TCP small queue contributes to 50% of overall performance, which suggests adapting the TCP/IP stack runtime to the upcoming resource disaggregated application is necessary. SN loads the tail of TxQs of every CN in sequence so that the number of CXL-like transactions is reduced, in return, the throughput is dropped by 25%. This emphasizes a prediction mechanism is needed to trade the SN’s load frequency for high throughput. The DRAM Cache plays the most significant role in the DFabric’s performance, as the latency of the CXL-attached memory pool is at most 10x of the local memory. However, such latency gap will be minimized if commercial CXL memory devices provide promised 170-250ns latency (ASPLOS23TPP, ).
Figure 13. The classification of the transactions from $CN_{0}$ (up) and $CN_{1}$ (down) within the DFabric in a 100ms time slice. We tag the time interval when the CN mainly receives packets from the SN and load/store data from/to the memory pool with purple and green shadows respectively.
In DFabric, some CNs of the same rack can temporarily use all the bandwidth of the NIC pool to minimize its communication period. To present such benefit, we allocate NIC pool’s receiving buffer from the uncacheable Region and capture the CXL-like memory transaction rate of CNs, while measuring the receiving bandwidth simultaneously. As shown in Figure 13. we find that two CNs utilize the NIC pool in a time-sharing manner. The period for computation, i.e., a CN load/stores the memory during its computation, is more than the one for communication, i.e., the other CN is receiving network traffic with the NIC pool. The peak memory bandwidth of the memory pool required by the NIC pool to receive network traffic is 2.9x of that accessed by CNs. This means the NIC pool requires a higher memory bandwidth of the memory pool to receive data, and the CN accessing the data with CXL needs a relatively low memory bandwidth, which can be satisfied by the CXL link connecting a CN with the memory pool. It proves the necessity of the co-existence of the memory pool and the NIC pool within the DFabric.
7. Discussion
Separating data and control plane in the SN is a common optimization to assign tasks to the appropriate computing resources (Guo22Clio, ). DFabric assigns the LPPU with the control plane, which is not often evoked, complicated, full of control paths, and hard to accelerate. DFabric leaves the data plane to the dedicated ASIC, which provides high data processing throughput. However, concerning the polling tasks, the multi-core CPU or smartNIC may have enough capability (Chen23Cowbird, ). For example, we can bind one core to poll one CN or NIC.
In DFabric, the sub-flow is mapped to one NIC based on its physical_queue depth. However, the factors, such as congestion and hops in the core network, should also be considered. The commercial data centers’ topologies are required to simulate various congestion situations. We hypothesize no congestion and one hop in this paper and leave them for future work.
8. Related Work
Prior works leverage multi-port NICs or add additional NICs to expand host egress capacity (NSDI15GRIN, ), which is effective given the huge gap between interconnects and networks. GPU clusters such as ELUPS (Micro14Zang, ) and EFLOPS (HPCA20EFLOPS, ) expand the rack-level bandwidth by bounding each GPU to a single NIC to avoid PCIe contention. Disaggregated Rack Architecture (DRA) (Micro14Zang, ; ISCA13Ladon, ; ANCS14Marlin, ) uses PCIe interconnect to allow a host to use multiple NICs of the rack, which is configured in advance. Similar to ToR-based racks, the host in DRA can be limited by Integrated I/O controller.
9. Conclusion
We present a novel interconnect architecture, DFabric, that bridges the bandwidth gap between interconnects and networks with a NIC pool, and optimize its efficiency with several novel designs such as a shared memory pool and DRAM Cache. We build a dual-rack prototype and validate its performance with both microbenchmarks and real experiments.